OpenStudy (anonymous):

complete the square 6v^2+4v+11=0 how do i do this???

OpenStudy (anonymous):

6v^2+4v+11=0 6v^2+4v=-11 6(v^2+ 2/3 v )=-11 6(v^2+ 2/3 v + 1/9 - 1/9)=-11 6(v^2+ 2/3 v + 1/9 ) + 6/9=-11 so you can already see that the number you are going to have to add to both sides to get rid of -6/9, is by smaller (or I should say, has an absolute value that is by far smaller) than -11. this will result in imaginary values then... understand? if not, ask....

OpenStudy (anonymous):

in my last line, on the outside of the parenthesis , it should be -6/9 not +6/9

OpenStudy (anonymous):

Like this, 6(v^2+ 2/3 v + 1/9 ) - 6/9=-11